Complete combustion of 2.70 g of a hydrocarbon produced 8.68 g of CO2 and 2.96 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.
Answer #1

let in compound number of moles of C and H be x and y respectively

Number of moles of CO2 = mass of CO2 / molar mass CO2

= 8.68/44

= 0.1973

Number of moles of H2O = mass of H2O / molar mass H2O

= 2.96/18

= 0.1644

Since 1 mol of CO2 has 1 mol of C

Number of moles of C in CO2= 0.1973

so, x = 0.1973

Since 1 mol of H2O has 2 mol of H

Number of moles of H = 2*0.1644 = 0.3289

Divide by smallest:

C: 0.1973/0.1973 = 1

H: 0.3289/0.1973 = 1.67 = 5/3

Multiply by 3 to get simplest whole number ratio:

C: 1*3 = 3

H: 5/3 * 3 = 5

So empirical formula is:C3H5

Answer: C3H5